OpenStudy (anonymous):

Help me please? Please write the Order of Operations. There were 2 ways we reviewed to help you remember. Please write what each step is asking you to do.

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

Please Excuse My Dear Aunt Sally

OpenStudy (sleepyhead314):

but I'm old, they must've come up with a newer phrase by now xD

OpenStudy (anonymous):

My teacher still uses that

OpenStudy (sleepyhead314):

PEMDAS

OpenStudy (anonymous):

i cant remember what it means tho this year all my classes are hard and everything is thrown at me to fast and i dont get enough help (I feel kinda sad i have to use openstudy to get help idky;-;)

OpenStudy (sleepyhead314):

That's ok ^_^ it took me a long while to remember this too xD P = Parenthesis - do everything inside the parenthesis first E = exponents - [perhaps you haven't learned these yet] M = Multiplication D = Division -- do these next, but M/D can be interchangable -- do them left to right example = 4 / 2 * 3 = 2 * 3 = 6 A = Addition S = subtraction -- same left to right idea as M/D, but after you do both M and D

OpenStudy (anonymous):

Okay but im still confused how to write the Order of Operations e.e im really lost

OpenStudy (sleepyhead314):

What I just decribed above is the Order of Operations (Parenthsis, Exponents, Multiplication, Division, Addtion, Subtraction) and next to them is a discription of what it does/what order specifically

OpenStudy (anonymous):

Oh okay thank you i didn't realize it

OpenStudy (sleepyhead314):

that's alright ^_^ I understand that it's a lot of info sorry if it was a little confusing and feel free to ask me about anything you might be still confused on :)

OpenStudy (anonymous):

Alright thank you very much for your help :)

OpenStudy (anonymous):

@sleepyhead314 do you know if -10 is an integer? thats the last thing i swear :o

OpenStudy (sleepyhead314):

It is because it Integers Include Negative numbers ^_^

OpenStudy (sleepyhead314):

you can keep on asking :P it's not like I've got aynthing better to do xD xD

OpenStudy (anonymous):

Thank you xD

OpenStudy (anonymous):

can you help me Solve for the variable? 52(n) = 312

OpenStudy (sleepyhead314):

see how the 52 is attached to the "n" by multiplication? so if we want to get "n" alone, we have to Divide both sides by 52 to undo the multiplication

OpenStudy (sleepyhead314):

so it would become n = 312 / 52

OpenStudy (sleepyhead314):

because 52 / 52 [on the left] = 1

OpenStudy (sleepyhead314):

and 1(n) = n

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

my friend is coming over in abit to study with me if i was gonna go over this with her how would i explain it?

OpenStudy (sleepyhead314):

To solve for a variable, you want to see what is stopping the variable from being alone and then try to undo it so if 4 + n = 10 ^^ the adding a 4 is stopping the "n" from being alone and to undo the 4, you would subtract 4 from both sides

OpenStudy (sleepyhead314):

does that make sense?

OpenStudy (anonymous):

Yeah. Did i ask how to find the variable for 52(n) = 312? im not sure if i did yet

OpenStudy (sleepyhead314):

yes you did ^_^ you can scroll up :)

OpenStudy (anonymous):

Okay im almost done i forgot i had something else do you think you can help me with one more thing it shouldnt take long?

OpenStudy (sleepyhead314):

of course ^_^

OpenStudy (anonymous):

Expression Variable Numerical Expression Equation Absolute Value Inequality Variable Expression

OpenStudy (anonymous):

just something else i gotta do but i dont understand it

OpenStudy (anonymous):

whats going on its writing everything down i put earlier

OpenStudy (sleepyhead314):

hmm what are you supposed to do with this? :)

OpenStudy (anonymous):

my account is messing up?

OpenStudy (anonymous):

my acc is messing up can we message each other? Cause its not working everything i put it translate to something i put down earlier

OpenStudy (sleepyhead314):

message how? I don't have my phone on me .-.

OpenStudy (anonymous):

Theres a little mail thingy on here its next to fanned users and notifications

OpenStudy (sleepyhead314):

right, sure, we can do that :)

OpenStudy (anonymous):

thank you

jigglypuff314 (jigglypuff314):

Sorry, sleepyhead has gone offline

OpenStudy (anonymous):

oh okay thank you for letting me know.